OpenStudy (ghdoru11):

Find the Fourier expansion for: f:R->R periodic T=2pi , f(x)=3x, x is between [-pi,pi]

OpenStudy (beginnersmind):

Why was this question closed?

OpenStudy (ghdoru11):

no one solve this and i have another question

OpenStudy (beginnersmind):

Anyway you can just go back to the definition of fourier coefficients: \[a_n = \int_{-\pi}^{\pi} 3x*sin(nx)dx\] \[b_n = \int_{-\pi}^{\pi} 3x*cos(nx)dx\] then evaluate the integrals. Use the substitution t = nx and integration by parts with u being the linear term and v' the trigonometeric part.

OpenStudy (irishboy123):

yeah, and it's odd so you ignore the cosine term

OpenStudy (beginnersmind):

Right, that saves a lot of work :)

OpenStudy (ghdoru11):

so just the first integral is neded?

OpenStudy (beginnersmind):

Yes, \(b_n\) = 0 for every n.

OpenStudy (ghdoru11):

thank you so much guys